### v1

# Automated Defects Detection and Fix in Logging Statement

Abstract.
Developers write logging statements to monitor software by capturing runtime behaviors, but misleading logs can even complicate software maintenance by obscuring real software activities. Existing research on analyzing quality issues of logging statements is limited, primarily focusing on detecting a singular type of defect and relying on manual intervention for fixes.
To address the limitation, we initiate a pilot study that pinpoints four distinct types of defects in logging statements through the analysis of real-world log-centric changes. We then propose LogFixer, a two-stage framework for automatically detecting and updating logging statements. In the offline stage, LogFixer constructs a similarity-based classifier on a set of synthetic defective logging statements to identify specific defect types. During the online testing phase, this classifier first evaluates logging statements in a given code snippet to determine the necessity and type of improvements required. Then, LogFixer constructs type-aware prompts from historical logging update changes for an LLM-based recommendation framework to suggest updates on these statements.
We evaluate the effectiveness of LogFixer on a dataset containing real-world logging changes, a synthetic dataset, and new real-world projects without having identified any defect previously. The results indicate that our approach is highly effective in detecting logging defects, achieving an F1 score of 0.625. Additionally, it exhibits significant improvements in suggesting precise static text and dynamic variables, with enhancements of 48.12% and 24.90%, respectively. Furthermore, LogFixer achieves a 61.49% success rate in recommending correct updates on new real-world projects. We reported 40 problematic logging statements and their fixes to GitHub via pull requests, resulting in 25 changes confirmed and merged across 11 different projects.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Logging is a fundamental practice to provide visibility of systems status, especially for revealing complicated software activities in the production environments (He et al., 2021; Jiang et al., 2024a). Listing 1 shows a logging statement in if-block containing three ingredients: logging level (i.e., debug), static contents (i.e., Left sub-range fully inconsistent), and dynamic variables (i.e., depth and right(Jiang et al., 2024b; Huang et al., 2024).
Log messages are generated when the logging statements are executed, which serve as primary resources to monitor system run-time behavior (e.g., detecting anomalies (Liu et al., 2023; Yu et al., 2024; Zhang et al., 2024b) and analyzing root causes (Li et al., 2024b; Chen et al., 2024b)). Consequently, the quality of logging statements becomes essential in offering accurate system information to engineers. Imprecise or misleading logging statements can give engineers incorrect information, leading to longer error diagnosis times and delaying their response to anomalies (Yang et al., 2024). Taking the above Cassandra code in Listing 1 as an example [^1] https://issues.apache.org/jira/browse/CASSANDRA-15173, the static content of this logging statement is inconsistent with the dynamic variable: the static content indicates the left tree is recorded, but the dynamic variable actually represents the right tree. When executed, the corresponding produced log message will mislead engineers into looking the wrong tree.
Listing 1: A logging statement example from Cassandra.
static int differenceHelper(...) { ...
else if (!lreso) {
log.debug("({}) Left sub-range fully inconsistent {}", depth, right);
ldiff = FULLY_INCONSISTENT;
Despite the prevalence of logging in modern software, the quality issue in logging statements receives insufficient attention due to the following challenges (Gholamian and Ward, 2021). Firstly, the large volume of logging statements within a codebase makes manual inspection impractical (Yuan et al., 2012b; Chen et al., 2024a). Secondly, because defects in logging statements do not directly affect program execution, the functional testing suites fail to detect such issues (Bouzenia and Pradel, 2023). Consequently, logging statements written by individual developers may not undergo a thorough review to examine their quality before deployment (Zhu et al., 2015).
While most studies have developed tools to decide logging locations (Zhao et al., 2017; Li et al., 2020; Cândido et al., 2021) and suggest logging content (Ding et al., 2022; Liu et al., 2019; Mizouchi et al., 2019; Li et al., 2024a; Xu et al., 2024; Li et al., 2021a), only a few have looked into identifying unqualified logging statements in existing software projects (Ding et al., 2023; Bouzenia and Pradel, 2023; Chen and Jiang, 2017). For instance, Ding et al. (Ding et al., 2023) studied the temporal relation between logging statements and the actual code sequence, proposing a detection framework for temporal issues. Li et al. (Li et al., 2023a) analyzed the readability of logging statements and attempted an automated readability issue detection process. However, existing works suffer from two drawbacks: (1) Limited scope of analysis: Previous studies investigate the quality of logging statements from single perspectives (e.g., temporal defects, readability issues), failing to provide a comprehensive overview of logging quality issues. This results in limited practicality when applied to real-world applications encompassing multiple categories of defects. (2) Detection without repairing: Existing works solely focused on detecting certain defects in logging statements without offering suggestions for repairs. Therefore, these approaches require extra human intervention with domain-specific knowledge to correct the defects. Automating the process of updating logging statements remains an unexplored area.
In this paper, we begin with a pilot study to uncover the potential defects that can affect the quality of logging statements by analyzing the existing logging-related software updates. Upon analysis, we derive the following insights based on challenges in updating defective logging statements: (1) four objective defects affect the quality of logging statements variously, necessitating precise identification and tailored correction for each type; (2) although defective logging statements are common in logging-related updates, they constitute only a small percentage of all logging statements. This imbalanced distribution of data highlights the need to consider the resource overhead of our approach, particularly when employing large language models (LLM) for fixing; and (3) recent research shows that learning-based methods can detect defects in certain categories of logging statements (Bouzenia and Pradel, 2023; Ding et al., 2023). Additionally, repairing defects with contextual defect information yields better results compared to arbitrary fixing (Yang et al., 2023). Therefore, an effective strategy involves designing a defect detector, followed by suggesting repair updates based on the detection results.
Based on the above insights, we propose LogFixer, a framework for automatically detecting and updating defective logging statements. LogFixer contains an offline fine-tuning phase and an online phase for updating based upon detection. The offline part aims to develop a defective logging statement detector. We fine-tune a similarity-based classifier for defect type identification on a set of synthetic defective logging statements. During the online phase, LogFixer first applies the above fine-tuned classifier to determine whether the input code snippet needs improvement and identify the specific types of defects. Finally, LogFixer constructs defect type-aware prompts and incorporates surrounding code contexts for double-checking and suggesting logging statement updates.
We extensively evaluate LogFixer in three datasets, including an existing log patches dataset collected from real-world software history, a synthetic dataset containing artificial defects, and a new dataset with unidentified defects for human evaluation. Our approach achieves promising detection ability (0.625 at F1) and a significant boost of 48.12% and 24.90% in static text and dynamic variables update ability, respectively. LogFixer also demonstrates effective detection and updating capabilities on new project data, achieving an overall successful rate of 61.49%. To date, 25 out of 40 changes submitted to the GitHub project developers have been merged, underlining the practicality of LogFixer.
In summary, this paper makes the following contributions:
  • We conduct a pilot study and identify four types of common defects that developers are concerned about, which can affect the quality of logging statements.
  • We propose a two-phase framework to automatically detect and update potential defective logging statements. To the best of our knowledge, this is the first work to automatically fix defective logging statements.
  • Experiment results on three datasets demonstrate the effectiveness and practicality of LogFixer in repairing defective logging statements.
Paper Organization Section 2 conducts a pilot study to investigate the defects that affect the quality of logging statements. Section 3 introduces our framework for automating detecting and updating defective logging statements. Section 4 shows the implementation details. Section 5 describes the experimental results. Section 6 presents a discussion on the threats to validity. Section 7 discusses the related work. Section 8 concludes the paper.
2. Pilot Study
This section systematically analyzes defects that may compromise the quality of logging statements in real-world applications. Given that normal logging statements modifications occur concurrently with functional updates (Yuan et al., 2012a), we assume that the isolated commit changes to these statements are related to repairing quality issues (Chen and Jiang, 2017). Therefore, we focus on examining the log-centric change (i.e., LCC), which is the software commits that solely involve the modification (exclude addition and deletion) of one or more logging statements.
2.1. Dataset Construction
Most existing studies (Ding et al., 2023; Chen and Jiang, 2019) concentrate on only several (i.e., four to six) mature repositories (e.g., Hadoop and Zookeeper), which lacks a comprehensive overview of common logging quality issues in other nascent software. Instead, our pilot study is conducted based on a wider range of representative repositories, denoted as target repositories.
2.1.1. Select target repositories
We first collect projects from popular and developing Java repositories on Github, denoted as target repositories. Specifically, the repositories must satisfy the following criteria:
  • The number of commits for the projects should range between 1,000 and 100,000, ensuring sufficient community interest and maintenance.
  • Projects must have been created before December 31, 2019, and the last commit must be later than January 1, 2023. This criterion ensures that the projects have a significant history and sustained maintenance.
  • The number of stars for the projects must exceed 1,000, indicating the popularity of the projects.
We utilized the tool provided by (Dabic et al., 2021) to filter GitHub repositories based on the aforementioned criteria. After excluding forks and the repositories without logging statements, we obtained a total of 641 repositories as the target repositories for manual analysis. Table 1 shows statistics of these repositories.
Table 1. Statistics of the studied repositories.
Avg # Code line Fork Commit Issue
Target repositories 454,157 1,886 7,618 1,772
2.1.2. Extracting log-centric changes
We then track changes in logging statements based on the evolution history of the target repository. In specific, we start with developing scripts to execute the git log command on the historical codebase from target repositories. These scripts extract the source code and relevant metadata for each commit, including identifiers such as the commit hash and content. Then, we employ ChangeDistiller (Fluri et al., 2007) to localize modifications between consecutive versions. Furthermore, considering that logging statements typically describe the code context within a method (Fu et al., 2014), we isolate functions containing logging statements from the repository. This aligns with previous intra-function logging practices research settings (Yuan et al., 2012a; Shang et al., 2015). Next, we filter out the commits including functional code changes, reserving those where all modifications are related to logging lines. Ultimately, we identified 10,137 LCCs.
2.2. Manual Study Design
To conduct manual analysis, we follow the sampling procedure in previous studies (Chen et al., 2018; Li et al., 2023b). Specifically, we employ stratified sampling (Pirzadeh et al., 2011) to select 500 LCC instances with a 95% confidence level and 5% confidence interval (Boslaugh, 2012). Considering the complexity of this investigation, we enlist two experienced co-authors as annotators, each with over seven years of programming experience. Given each LCC with its associated method, they are asked to investigate the reason for changes by analyzing the data flow and structural context surrounding each logging statement. Based on this analysis, they identify the types of changes made by the programmers. The results in this phase have a Cohen’s Kappa (Sim and Wright, 2005) of 0.78, which is a substantial level of agreement.
Table 2. Summarized rationals for log-centric changes.
Type Number Percentage
Correct statement-code inconsis. 73 14.6%
Correct static-dynamic inconsis. 45 9.0%
Correct temporal inconsis. 11 2.2%
Correct typos/capitalization 129 25.8%
Add/delete statement info 105 21.0%
Change log level 124 24.8%
Refactor output format 8 1.6%
Others 5 1.0%
Total 500 100%
2.3. Defects Type in Log-centric Changes
Table 2 displays the manual analysis results, which summarizes seven reasons for updating logging statements. Our study centers on the first four categories because they relate to factual defects, result in objective description errors on system behavior and directly impact the quality of logging statements. Instead, the rest modifications (i.e., add/delete statement info change log level, and refactor output format) are subjective and can vary significantly according to the repository and developer. Notably, we observe that 51.6% (258/500) of the changes relate to the first four factual defects, highlighting the significant attention developers devote to these issues.
More specifically, we detail the symptoms and explain how each defect weakens the quality of logging statements as follows. Figure 1 offers examples of each defect from four different repositories.
2.3.1. Statement-code inconsistency
This defect indicates that the semantics of the logging statement are inconsistent with the surrounding code. The fundamental purpose of logging practice is to capture the runtime system activities of the current program for operation. A mismatch between the content of the logging statements and their code context leads to inconsistencies that can mislead engineers in failure diagnosis. For example, in Fig. 1(a), the DDL success event was mistakenly logged as Catalog update success.
2.3.2. Static-dynamic inconsistency
This refers to the discrepancy between static logging text and dynamic variables. Static text within logging statements must carry the meaning and implications of the associated dynamic variables. Otherwise, the logged variable strings are meaningless and hard to comprehend for log inspectors. As in Fig. 1(b), the static text “Current mDatasetCounter: ” incorrectly describes the dynamic variable mInodeCounter.get(). This misalignment could lead to confusion when debugging or reviewing logs.
(a) Statement code inconsistency at DdlThread.java in voltdb.
(b) Static dynamic inconsistency at MasterInfo.java in alluxio.
(c) Temporal relation inconsistency at AbstractResource.java in strimzi-kafka-operator.
(d) Readability issue at QuorumPeerConfig.java in zookeeper.
Figure 1. Real-world examples for four defects.
2.3.3. Temporal inconsistency
The tense of the logging text is inconsistent with the actual situation. The tense used in logging statements should align with the temporal context of the events. This adherence ensures clarity for tracking system events. In Fig. 1(c), the original logging statement uses the present continuous tense “Creating”, which implies that the action of creating the exec probe is still ongoing. However, the logging statement is placed immediately before the return statement, indicating that the creation of the probe has been completed by the time of the logging statement execution.
Figure 2. The overview framework of LogFixer.
2.3.4. Readability issues
The logging statement with spelling errors or abusive capitalization can undermine their readability. Eliminating these errors is crucial to prevent misunderstandings and ensuring that logs serve as clear and effective communication tools for developers and operators. Taking Fig. 1(d) as an example, ”Non-optimial” is a typo error.
Summary: The study uncovers four categories of factual defects (accounting for 51.6% LCC) that can affect the quality of logging statements.
3. Framework
3.1. Overview
[…]
To this end, we propose LogFixer, a framework for automatically detecting and updating defective logging statements. Fig. 2 illustrates the high-level view of LogFixer, which operates in two stages: offline fine-tuning and online updating. During the offline stage, the goal is to develop a defective logging statement detector. We first synthesize the defective logging statements with corpus collected from well-maintained repositories, ensuring high-quality code and maintenance standards. Then we use the synthetic data to fine-tune a similarity-based logging defect detector. The online phase involves two steps: detecting defective logging statements and updating them. We first use the fine-tuned detector to determine the $type_{s}$. Based on the defect information, we apply the analyzed knowledge into the LLM-based updater for logging statement update.
3.2. Defective Logging Statement Synthesis
To prepare the corpus for fine-tuning the detector, the initial step involves synthesizing negative samples. These samples are sourced from well-maintained repositories and mutated by text mutation strategies, as the corpus for tuning defective logging statement detectors. This foundational step is crucial as the training data quality employed in developing the model significantly influences its overall performance. Previous studies (Bouzenia and Pradel, 2023) emphasize three crucial objectives in the data generation methodology: realism, diversity, and scalability. Realism ensures that the dataset mirrors actual errors that developers might encounter. Diversity is vital for enabling the model to recognize various types of issues. Lastly, scalability is necessary since training a robust model typically demands a large dataset. In alignment with these objectives, we develop three techniques to synthesize defective logging statements for different defects, detailed as follows.
3.2.1. Mutation of words
For readability issues, we synthesize defective logging statements by making straightforward modifications to one word in the static constant part of the logging statement. Specifically, for simulating typos, we randomly select a word and alter its spelling based on the known typo dataset (Hagiwara and Mita, 2020; Microsoft Corporation, 2016). If the chosen word is not in the dataset, we then randomly add, delete, or change one of its characters. For issues with capitalization abuse, we randomly select a word and convert it to uppercase. We ensure that only one type of word mutation (e.g., executor $\rightarrow$ executer) is applied to each logging statement in modification.
public void stop() {
for (InstanceExecutor executor : Executors){
executor.stop();}
LOG.info("To stop Stream executer");
// *** executor -> executer ***
streamExecutor.stop();
3.2.2. Mutation of tense
Inspired by previous work (Ding et al., 2023), for temporal inconsistency, we use spaCy (Honnibal et al., 2020), which is an open-source NLP library, to alter the tense of logging statements. Specifically, we employ dependency parsing (Zmigrod et al., 2021) and part-of-speech (POS) tagging (Coleman, 2005) to identify the main verbs in the logging statements. We consider the verb tagged as a head token in the dependency tree as the main verb. Finally, we change the main verbs to random tenses to mutate examples where logging statements do not align with the system events. For instance, we mutate starting to started before the sequence generator source starts in the following examples.
protected void doStart() throws FlumeException {
logger.info("Sequence generator source do started");
// *** starting -> started ***
sourceCounter.start(); }
3.2.3. LLM-based mutation
For defects requiring semantic mutation (statement-code inconsistency and static-dynamic inconsistency), using rule-based or heuristic methods can be challenging. Therefore, we employ LLM-based techniques to synthesize defective logging statements. These LLMs, trained on vast amounts of code, have been shown to handle code-related downstream tasks effectively (Zhang et al., 2024a; Xia et al., 2023; Yang et al., 2023). We use GPT-4o (OpenAI, 2024) to mutate the semantics for static content or dynamic variables. Specifically, for statement-code inconsistency, we prompt the LLM to carefully analyze the surrounding code and then alter the static content of logging statements to create semantic inconsistencies (e.g., recording opened while the actual event is close). For static-dynamic inconsistency, we instruct the LLM to randomly mutate the static content or dynamic variables to generate potential mismatches (e.g., recording exc while a directory path dir is expected).
if (channelRef.get() != null) {
channelRef.get().close();
LOG.debug("channel {} opened", remoteAddr);
// *** closed -> opened *** }
public postVisitDirectory(Path dir, Exception exc) {
LOGGER.info("> Deleting folder {}", exc);
// *** dir -> exc ***
Files.delete(dir); }
After generating defective logging statements using above techniques, we filter out the identical samples. This step is crucial as various data mutation techniques may inadvertently generate identical data. The presence of such duplicate data could potentially skew our detector’s learning process, leading to an overemphasis on homogenous data features.
3.3. Logging Statement Defects Detection
Considering that only a small portion of logging statements require update, and scanning all methods in a repository with LLM is resource-intensive, LogFixer employs a similarity-based classifier as the detector. This detector identifies defective logging statements and determines the types of defects. The introduction of the fine-tuned detector significantly reduces costs and enables subsequent LLM components to update the logging statements in a more efficient and targeted manner.
Specifically, this phase is responsible for handling a multi-class classification task. Given a logging statement $s_{original}$ with its associated context (i.e., code snippet), the goal is to determine whether $s_{original}$ has a specific type of defect. We resolve the problem through the following steps:
3.3.1. Tokenization and Embedding
This step converts the input code snippet into a series of tokens. In this study, we use UniXcoder (Guo et al., 2022), which is a unified cross-modal pre-trained model for programming languages that are based on a multi-layer transformer and utilizes mask attention matrices (Dong et al., 2019) with prefix adapters to control the access to context for each token. We adapt the pre-trained tokenizer provided by the Unixcoder and embed the series of tokens into fixed-size vectors. The fine-tuning process is end-to-end where the parameters of the encoder layers are adjusted to enhance the representation vectors, allowing these vectors to better adapt to detection task.
3.3.2. Similarity-based multiple classification
After obtaining the corresponding representations for the target method and logging statement, we then devise a multi-class classifier to identify defects. Intuitively, a “good” logging statement should closely align with the semantic meaning of its adjacent code. Hence, we have developed a similarity-based classification model instead of applying a basic multi-classifier with fully connected layers. Specifically, the classifier concatenates the embedding of the logging statement $l$, and its corresponding code snippet $s$. Here are several benefits of doing so. First, explicitly identifying the target logging statements allows the model to focus more on the semantic meaning related to target logging statements. Second, it addresses the issue of multiple logging statements within a single method, as specifying $l$ informs the model which statement to analyze. The loss function, shown in Equation 1, includes a cross-entropy loss and a similarity-based term. The first part is the standard cross-entropy loss (Rumelhart et al., 1986), commonly applied in machine learning and optimization areas. The second part is a similarity-based loss term, which forces the target logging statement vectors to be similar to the surrounding code snippet vectors by maximizing their cosine similarity. $\alpha$ is a weighting parameter to control the influence of each term. The last term is added to ensure the total loss function remains above zero.
(1) $loss=-(\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(p_{i,c})+\alpha\frac{1}{N}\sum_{i=1}^{N}\frac{l_{i}\cdot s_{i}}{\|l_{i}\|\|s_{i}\|}-\alpha)$
In the equation, $C$, $N$ represent the total number of defect types and samples, respectively. The variable $y_{i,c}$ is an indicator for the ground truth label for the $c$-th type of the $i$-th sample, while $p_{i,c}$ donates the predicted probability for the same type. The vector $l_{i}$ and $s_{i}$ correspond to the vector representations of the target logging statement and the surrounding code snippet for the $i$-th sample, respectively.
3.4. Defective Logging Statement Update
Figure 3. The recommendation framework on a static-dynamic inconsistency case.
Following the defect classification phase, where potential defective logging statements are identified, we proceed with a multi-expert collaboration framework to repair these statements, inspired by the cognitive synergy (Goertzel, 2009). In particular, we employ two LLM-based experts: a checker and an updater. Fig. 3 depicts a synergy between two experts, as well as the specific duty, and the prompt for each expert. The checker’s firstly validates the classification results from the detector, and then employs LLMs to provide supplementary semantic information. Afterward, the updater is responsible for implementing logging statement updates according to the advice provided by the checker.
3.4.1. Checker
The role of checker is to reduce the false positives brought by the similarity-based detector, where non-defective logging statements could be mistakenly marked as requiring update due to the extremely imbalanced data in real-world situations. To address this issue, we have designed a checker with two objectives. The first objective is to determine whether the detector’s decisions on updating logging statements are necessary, aiming to eliminate false positives. The second objective is to leverage the extensive knowledge of the LLM to offer additional semantic insights, thereby enhancing the effectiveness of subsequent experts.
Specifically, we use the information generated by the detector (i.e., target logging statement, surrounding code, defect type) as the input of the checker. Ultimately, the checker is expected to generate its judgment, rationale, and semantic information of the logging statements and the surrounding code. This information will be passed on to the next expert for reference.
3.4.2. Updater
Upon checking, we identified the target logging statements and their corresponding defect types. The next step involves employing the updater to modify these logging statements.
We utilize log-centric changes as supplementary knowledge to guide the LLM in learning the past update patterns from maintainers. We apply a defect type-aware strategy to select proper historical LCC examples for the updater’s reference. In specific, for defect types sensitive to project characteristics (e.g., temporal relation inconsistency and readability issues), we exclusively select the LCCs from the same project to maintain the consistency in the project’s logging style (Li et al., 2024a). In contrast, for defects related to semantic information inconsistency that are unaffected by logging style and inter-project variations, we incorporate LCCs from multiple projects for the updater. We employ the BM25 (Robertson et al., 2009) ranking algorithm to identify LCCs that are the most similar to the target logging statements.
4. Implementation
4.1. Fine-tuning Dataset
To construct high-quality data for fine-tuning the detector, we apply the technique mentioned in 3.2 to generate 5K defective instances for each type. Meanwhile, we consider logging statements from well-maintained repositories as defect-free because they adhere to stringent community guidelines and thorough maintenance protocols. Specifically, we select well-maintained repositories from the target repositories based on the following criteria:
  • The initiators of the repository must be companies, such as those from Google, Apache, Facebook, etc. This ensures that the projects are backed by experienced teams leading to better development practices and higher-quality code.
  • The repository must allow for the submission of issues; open community discussions are beneficial for developers to maintain the project.
  • Projects must have a license, ensuring the availability of the code for widespread use.
We ultimately considered a total of 161 projects, encompassing 128K methods with logging statements, as well-maintained repositories. In addition to 20K defect instances, we also randomly selected 20K examples from the well-maintained repositories as positive samples, leading to 40K samples in total. We split the data into training, validation, and test sets with a ratio of 8:1:1 using stratified sampling (Pirzadeh et al., 2011). This results in a total of 32K samples for fine-tuning the detector.
4.2. Implementation Details
During LLM-based defective logging statement synthesis, we use the GPT-4o (OpenAI, 2024) API provided by Openai as the default LLM model. We use UniXcoder-base as the tokenization and embedding tool. We use Adam (Kingma and Ba, 2017) as the optimizer and set the learning rate to 5e-5, Adam epsilon to 1e-8, and dropout rate to 0.1 when fine-tuning the similarity-based classification model. We fine-tune the model for 10 epochs and set the max token number to 1024. We run them on a server with Dual Intel(R) Gold 6326 CPU, 512GB physical memory, and 4x NVIDIA RTX 3090 GPU. The OS version is Ubuntu 18.04.
5. Evaluation
We evaluate LogFixer by answering the following research questions:
RQ1: How effective is LogFixer at detecting and enhancing defective logging statements?
RQ2: How do different components affect the performance of LogFixer?
RQ3: How generalizable is LogFixer for different LLM backbones?
RQ4: How is the practitioners’ feedback on LogFixer?
5.1. Evaluation Dataset
To evaluate the performance of LogFixer, we conducted experiments on three distinct datasets.
5.1.1. Existing logpatches
We utilize the data annotated in the manual study (2.2) as existing logpatches. Each sample forms a pair, consisting of an original defective logging statement and its corrected version, culminating in a total of 258 defective pairs. Additionally, we sample 258 non-defect logging statements from well-maintained repositories, which brings our total case count to 516.. Note that we exclude these data from the training set.
5.1.2. Synthetic test data
As mentioned by 4.1, we used a stratified sampled test set including 4,000 samples, as the synthetic test data.
5.1.3. New repositories
To evaluate the effectiveness of LogFixer in addressing previously unidentified defects, we randomly sampled 100K log-code pairs outside the collected well-maintained dataset for human evaluation. These samples were deliberately excluded from all previous processes, thereby guaranteeing that the evaluation was conducted on entirely fresh data.
5.2. Metrics
For the detection ability of LogFixer, we use Precision, Recall, and F1-macro to measure multiple classification performance. Formally, $Precision=\frac{TP}{TP+FP}$, $Recall=\frac{TP}{TP+FN}$, $F_{1}=2*(\frac{Precision*Recall}{Precision+Recall})$, and $F_{1}^{macro}=\frac{1}{N}\sum_{i=1}^{N}F_{1}^{i}$. We use $F_{1}^{macro}$ to give equal weight to each class, avoiding the issue of inflating the weight of the class with more instances.
For the updating ability, we use the same metrics as logging statement generation task. Different metrics are used to evaluate different parts of the logging statements.
(1) Static content. In alignment with recent research efforts (Li et al., 2023b; Mastropaolo et al., 2022), the quality of updated logging texts is evaluated using two established metrics from the field of machine translation: BLEU-K (where K = 1, 2, 4) and ROUGE-K (where K = 1, 2, L), to measure the overlap in terms of K-grams between the updated and actual logging texts.
(2) Dynamic variables. Following the recent work (Li et al., 2023b), we utilize special Precision, Recall, and F1 to assess the accuracy of updated logging variables For each logging statement, let $S_{ud}$ represent the variables in the LogFixer updates, and $S_{gt}$ denote the variables in the actual logging statement. We calculate and report the proportion of correctly updated variables ($precision$=$\frac{S_{ud}\cap S_{gt}}{S_{ud}}$), the proportion of actual variables that are correctly predicted by the model ($recall$=$\frac{S_{ud}\cap S_{gt}}{S_{gt}}$), and their harmonic mean ($F_{1}$=$2*\frac{Precision*Recall}{Precision+Recall}$).
The above metrics of updating evaluation, are widely used to assess the quality of text generation by calculating the similarity between generated log text and actual ones. However, a high similarity score in this task is not necessarily related to a significant improvement, as the defective logging statements prior to the update could already bear a close resemblance to the ground truth (i.e., the logging statements updated by developers). Consequently, we propose a novel metric named Improvement Coefficient (IC), which focuses on the change in these metrics before and after the updates. The formula 2 reflects the proportion of the improvement achieved after the updates relative to the maximum possible improvement. Specifically, $m\in\{BLEU,ROUGE,Precision,Recall,F_{1}\}$. $m_{origin}$ and $m_{update}$ denote the metrics before and after the updates, respectively.
[…]
5.3. Baseline
Table 3. The detection results on synthetic data and existing logpatches.
Precision Recall F1
Existing logpatches
CodeT5+ 0.314 0.062 0.096
LogFixer 0.725 0.552 0.625
Synthetic data
CodeT5+ 0.366 0.075 0.115
LogFixer 0.925 0.673 0.776
To provide a comprehensive evaluation, we compare the detection ability and update ability of LogFixer with representative methods, with details as below.
For the detection ability, we pre-trained a CodeT5+ (Wang et al., 2023) model based on the approach proposed by CMI-Finder (Bouzenia and Pradel, 2023) to serve as a baseline for comparison, which has shown good performance in the binary classification task of message-condition inconsistency.
For the updating ability, due to the lack of related research, we employed several different prompting methods as baselines. Specifically, we utilized instruction prompting (Instruct), in-context learning (ICL), and prompts with detection information (DET). Instruction prompting involves directly instructing the LLM to update defective logging statements without providing additional context. In ICL, we retrieve examples to help the model better understand the objective; we randomly sampled an add-delete pair from the LCC to create these examples. For DET, we provided the LLM relevant information about the type of defect and used prompts to guide the LLM to first explain why the target logging statement contains the specific defect before updating it.
5.4. RQ1: Effectiveness of LogFixer
Table 4. The updating result on existing logpatches and synthetic data.
{NiceTabular}
l—llllll—lll Static texts Dynamic variables
BLEU-1 BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L Precision Recall F1
Existing logpatches
GPT4o-Instruct 0.690 (4.90$\uparrow$) 0.587 (12.86$\uparrow$) 0.268 (0.94$\uparrow$) 0.746 (13.01$\uparrow$) 0.593 (17.44$\uparrow$) 0.745 (12.96$\uparrow$) 0.272 (0.13$\uparrow$) 0.256 (1.59$\uparrow$) 0.264 (1.15$\uparrow$)
GPT4o-ICL 0.681 (2.14$\uparrow$) 0.588 (13.08$\uparrow$) 0.262 (0.14$\uparrow$) 0.750 (14.38$\uparrow$) 0.603 (19.47$\uparrow$) 0.748 (13.99$\uparrow$) 0.300 (3.97$\uparrow$) 0.307 (21.82$\uparrow$) 0.304 (16.47$\uparrow$)
GPT4o-DET 0.701 (8.28$\uparrow$) 0.620 (19.83$\uparrow$) 0.327 (8.93$\uparrow$) 0.781 (25.00$\uparrow$) 0.650 (29.01$\uparrow$) 0.779 (24.57$\uparrow$) 0.315 (6.03$\uparrow$) 0.296 (17.46$\uparrow$) 0.305 (16.85$\uparrow$)
LogFixer 0.799 (38.34$\uparrow$) 0.721 (41.13$\uparrow$) 0.384 (16.64$\uparrow$) 0.849 (48.28$\uparrow$) 0.738 (46.85$\uparrow$) 0.848 (48.12$\uparrow$) 0.439 (23.04$\uparrow$) 0.451 (26.60$\uparrow$) 0.445 (24.90$\uparrow$)
Synthetic data
GPT4o-Instruct 0.546 (-7.32$\downarrow$) 0.465 (-7.86$\downarrow$) 0.316 (-2.54$\downarrow$) 0.634 (-6.08$\downarrow$) 0.499 (-9.15$\downarrow$) 0.632 (-6.35$\downarrow$) 0.696 (14.85$\uparrow$) 0.707 (42.20$\uparrow$) 0.702 (32.57$\uparrow$)
GPT4o-ICL 0.545 (-7.56$\downarrow$) 0.468 (-7.25$\downarrow$) 0.313 (-3.00$\downarrow$) 0.639 (-4.64$\downarrow$) 0.508 (-7.19$\downarrow$) 0.637 (-4.91$\downarrow$) 0.683 (11.20$\uparrow$) 0.664 (33.73$\uparrow$) 0.673 (26.02$\uparrow$)
GPT4o-DET 0.640 (14.89$\uparrow$) 0.579 (15.12$\uparrow$) 0.371 (5.69$\uparrow$) 0.715 (17.39$\uparrow$) 0.636 (20.69$\uparrow$) 0.714 (17.34$\uparrow$) 0.752 (30.53$\uparrow$) 0.761 (52.85$\uparrow$) 0.756 (44.79$\uparrow$)
LogFixer 0.659 (19.39$\uparrow$) 0.592 (17.74$\uparrow$) 0.436 (15.44$\uparrow$) 0.753 (28.40$\uparrow$) 0.648 (23.31$\uparrow$) 0.751 (28.03$\uparrow$) 0.795 (42.57$\uparrow$) 0.778 (56.21$\uparrow$) 0.787 (51.80$\uparrow$)
Approach. The effectiveness of LogFixer should be evaluated from two aspects: first, the detection capability, specifically the capability to identify defective logging statements; second, the updating capability, which pertains to the similarity between the revised logging statements (i.e., prediction) and the actual ground truth (i.e., repository update history).
Result. Table 3 shows the detection results with different classification models. Overall, our proposed detector significantly outperforms CodeT5+ across all metrics on both existing logpatches and synthetic datasets. This could be because the T5 architecture might struggle with complex code comprehension, as it lacks the utilization of structural features in the programming language (e.g., AST).
Table 5.4 illustrates the updating performance of static content. Compared to the three baselines, LogFixer demonstrates significant performance improvements across various metrics on two datasets. Specifically, for GPT4o-Instruct and GPT4o-ICL, which lack defect information, the improvement on the logpatches dataset is minimal (0.14$\uparrow$ IC on BLEU-4), and there is even a regression (3.00$\downarrow$) on synthetic data. However, GPT4o-DET, which includes defect information, shows a certain level of improvement, highlighting the importance of defect information.
Answer to RQ1: LogFixer demonstrates strong capabilities in defect detection and updating. It significantly outperforms CodeT5+ in detection with an F1 score of 0.625. Regarding its update ability, it improves static text and dynamic variables by 48.12% and 24.90%, respectively.
5.5. RQ2: Effects of Different Components
Approach. We consider the contribution of different components to LogFixer ’s performance from two perspectives.
For detection capability, we first experiment with several different configurations on PLM to understand how hyperparameters affect our detection results. Specifically, we change the epoch in fine-tuning and the $\alpha$ in loss function to control the importance of traditional cross-entropy loss and cosine similarity regularization term. Then we remove the checker LLM agent and use only the PLM for detection as a comparison. This will demonstrate the necessity of using the checker.
For updating capability, we compare the results with the scenario where the detection module is entirely removed and the LLM directly updates the logging statements that need to improve. This will illustrate the necessity of the detection module.
Result. For hyperparameters in PLM, the default is ten epochs and $\alpha$ is 0.5. Experimental results in figure 6 indicate that both reducing the training by 5 epochs (0.05$\downarrow$|↓ | 0.08$\downarrow$) or extending it by 5 epochs (0.03$\downarrow$|↓ | 0.04$\downarrow$) lead to a certain degree of performance degradation on both existing logpatches and synthetic data. Regarding the hyperparameter $\alpha$, setting it to 0 (using BCE loss directly) or setting it to 2 (giving higher weight to the similarity-based term) also results in significant performance degradation. This demonstrates the effectiveness of our designed loss function.
Figure 6 illustrates the detection performance of the framework with and without the checker. We observe that using the checker does not result in a significant change in the F1 score. Additionally, the precision on both datasets is significantly improved (0.15$\uparrow$|↑ | 0.11$\uparrow$), while the recall is notably reduced (0.20$\downarrow$|↓ | 0.04$\downarrow$). This aligns with our design rationale for the checker, which is to filter out the massive false positives generated by the similarity-based classifier.
Figure 6 illustrates the update performance of the framework with and without the detection component. We observe a significant decrease in various evaluation metrics across both datasets when the detection component is not used. Additionally, even without the detection component, the update component still provides some degree of repair (13.99$\uparrow$ IC on ROUGE-L) on the existing logpatches. However, its performance on the synthetic dataset is worse than not performing updates at all (10.98$\downarrow$ IC on ROUGE-L).
Figure 4. Hyperparameter analysis.
Figure 5. Detection results w/wo checker.
Figure 6. Updating results w/wo detection.
Answer to RQ2: Each component of LogFixer contributes variously to its overall performance. Adjustments to hyperparameters or removal of the checker or detection components result in performance decline.
5.6. RQ3: Generalizability of LogFixer
Approach. We evaluate the capability of LogFixer using code-specific (Phind-CodeLlama (Hugging Face, 2024), Deepseek-coder (Zhu et al., 2024)) and general (Claude3.5-sonnet (Anthropic, 2024), Llama3 (AI@Meta, 2024), GPT3.5 (OpenAI., 2022), GPT4o (OpenAI, 2024)) LLMs on both synthetic data and existing logpatches. This can illustrate the generalizability of our framework, which means that we don’t rely on the ability of specific LLM backbones.
Result. Table 5 shows the performance of LogFixer with six different backbone models. First, we observe that our framework can consistently perform well on most backbones in terms of all metrics by a large margin. On average, all models get 0.614 F1-macro for detection ability. From the perspective of update ability, models have an average of IC 13.19$\uparrow$ (reflected by BLEU-4), 38.68$\uparrow$ (reflected by ROUGE-L), and 12.52$\uparrow$ (reflected by F1) respectively.
The results demonstrate the generalizability of LogFixer across various backbone models. We anticipate that the performance of LogFixer can be further enhanced with advancements in LLMs.
Table 5. The performance of LogFixer with different backbone models.
Detection Updating text Updating variable
F1-macro BLEU-4 ROUGE-L F1
CodeLlama-34b-v2 0.613 0.377 (15.70$\uparrow$) 0.856 (50.85$\uparrow$) 0.319 (7.85$\uparrow$)
Deepseek-coder-v2 0.642 0.371 (14.88$\uparrow$) 0.818 (37.88$\uparrow$) 0.309 (6.50$\uparrow$)
Claude3.5-sonnet 0.637 0.355 (12.72$\uparrow$) 0.810 (36.15$\uparrow$) 0.359 (13.26$\uparrow$)
Llama3-70b 0.594 0.353 (12.45$\uparrow$) 0.819 (38.23$\uparrow$) 0.359 (13.26$\uparrow$)
GPT3.5 0.573 0.311 (6.77$\uparrow$) 0.771 (21.84$\uparrow$) 0.330 (9.34$\uparrow$)
GPT4o (original) 0.625 0.384 (16.64$\uparrow$) 0.848 (48.12$\uparrow$) 0.445 (24.90$\uparrow$)
Answer to RQ3: LogFixer showcases great generalization ability by consistently enhancing the quality of logging statements across various LLMs, including code-specific and general ones.
5.7. RQ4: Practioners’ feedback on LogFixer
In this RQ, we evaluate LogFixer based on the practitioners’ feedback from two prospective.
1) Human evaluation
Approach We aim to assess if LogFixer can accurately identify and update unseen defective logging statements while achieving a reasonable successful rate (SR). For this, we employ the new repositories dataset (Sec.  5.1.3), which has no prior identified logging statement defects and modifications. Two authors manually verified each case to confirm whether the detection is accurate and if the update is correct. If both annotators unanimously agree, we consider it a success case. Finally, we compute the success rate as the ratio of the number of successful cases to the total number of cases.
Result LogFixer reports 149 defective logging statements out of 100K data. The two authors achieve a Cohan’s Kappa of 0.84 and finally identify 91 successful detections and updates after reaching a consensus, resulting in a 61.49% SR. This outcome demonstrates the low false positive rate of LogFixer, enabling developers to conserve time in scrutinizing meaningless updates. Table 6 demonstrates SR for LogFixer across different types of defects, with the highest SR of 88.57% observed for readability issues, and the lowest at 46.15% for static-dynamic inconsistency.
2) Developer’s feedback
Approach. To understand the importance of fixing identified defects from the developers’ perspective, we randomly sample 40 updated logging cases and submit pull requests (PRs) to the repository maintainers. For each pull request, we include the defects type and the improved logging statements suggested by LogFixer. Developers’ willingness to invest effort in testing, validating, and accepting our PRs serves as a testament to the effectiveness of our approach and the perceived value of the proposed logging quality improvement.
Result. To date, 25 out of 40 proposed changes have been accepted and merged into the codebases, and an additional 3 cases have been recognized as bugs and are incorporated in future version. Table 7 shows five representative real-world defective logging statements updated by LogFixer (four successful cases and one failure case), with an in-depth discussion for each case in the “explanation” column.
Table 6. The SR for LogFixer in new data, ”SCI”, ”SDI”, ”TI”, and ”RI” indicate the statement-code, static-dynamic, temporal inconsistency, and readability issues, respectively.
Defect type SCI SDI TI RI
Successful rate 62.07% 46.15% 63.16% 88.57%
Table 7. The real-world defective logging statement update cases suggested by LogFixer.
Information Update code Explanation
Statement-code inconsistency merged by infinispan (Anonymous, 2024b) public CompletionStage<Void> rollbackAsync (){ - log.tracef("Transaction.commit() invoked in transaction with Xid=%s", xid); + log.tracef("Transaction.rollback() invoked in transaction with Xid=%s", xid); The older statement was misleading as it implies that a commit operation was undertaken, whereas the transaction actually involves rolling back.
Static-dynamic inconsistency merged by trino (Anonymous, 2024c) private internalCompilePageWithPositionComparator (...) { - log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortChannels); + log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortOrders); The original statement logged variable sortChannels twice, mistakenly applying it to record both the channels and orders in the message.
Temporal inconsistency merged by openhab-addons (Anonymous, 2024d) private class DigiplexReceiverThread extends Thread { - logger.debug("Receiver thread started"); + logger.debug("Starting receiver thread"); Optional<String> message = readLineBlocking(); The old statement used the past tense (“started”) suggesting that the thread had already started. However, in reality, the thread initiation occurs later in the code execution.
Readability issue merged by openhab-addons (Anonymous, 2024a) public PentairIntelliFloHandler(Thing thing) { ... - logger.debug("Intellflo received refresh command"); + logger.debug("IntelliFlo received refresh command"); } The original one contains typo Intellflo. LogFixer successfully update it to IntelliFlo (a pool pump) rather than Intelliflo (a software platform).
A failure case case offered from LogFixer (false positive). public Future<Integer> scaleDown(...) { while (next > scaleTo) { next--; - log.info("Scaling down from {} to {}", next + 1, next); + log.info("Scaling down from {} to {}", next, next - 1); LogFixer erroneously reports the logging statement as defective due to a misinterpretation of next+1 as representing the current number. This error suggests that LogFixer may struggle to calculate changes in numerical variables within programs, attributing to the limited mathematical ability of LLMs.
To save space, some parts of the update code snippets are abbreviated.
Answer to RQ4: In repositories containing previously unidentified defects, LogFixer suggests 61.49% precise logging fixes according to human evaluation, illustrating its real-world practicality. Out of 40 changes reported, developers have merged 25 cases, further validating its usefulness.
6. Threats to Validity
This section discuss the threats to validity of LogFixer.
Potential data leakage. We conduct experiments using various LLM backbones. Because the training data for these models are undisclosed, there is a risk of potential data leakage. To mitigate this, we initiate our data collection in April 2024, postdating the training of most LLMs. Furthermore, our methods also demonstrate strong performance on synthetic datasets, suggesting that the improvements in logging statements are from inference instead of memorization of training data.
Fine-tuning dataset construction. The fine-tuning dataset construction process may raise concerns. Firstly, similar to logging-related works (Mastropaolo et al., 2022; Li et al., 2020; Liu et al., 2022), the basic assumption is that the logging statements in the latest programs are defect-free. Secondly, due to the required fine-tuning set scale, manually labeling a large number of negative examples would be impractical. Consequently, we employed mutation techniques to automatically synthesize defective examples, potentially leading to the inclusion of suboptimal negative examples. To mitigate this, the selection of data source is limited to well-maintained codebases (details in 4.1), which are considered as positive data. Additionally, we devised various mutation methods to ensure the diversity, realism, and scalability of synthetic data.
The selection of programming languages. We source data from 641 Java projects, which may impact the generalizability of our method for other languages. Following previous study (Ding et al., 2023), we choose Java because it is a mature language with multiple built-in logging and third-party logging frameworks. Nevertheless, applying LogFixer to other programming languages may affect its performance because of distinct language features. While fine-tuning detectors on other language-specific datasets is a promising solution, we aim to incorporate more languages in future work.
Human involvement. We conduct manual studies to identify defects in logging statements and human evaluation of new repositories. This human-involved process may introduce bias during annotation. To minimize this threat, we engage two independent annotators with expertise in the field, and ensure that any discrepancies are resolved through discussion until a consensus is reached. Besides, we get 0.78 Cohen’s Kappa for manual defeat identification and 0.84 for human evaluation, which is a substantial level of agreement.
7. Related Work
Studies on maintaining high-quality logging statement. Many related works focused on potential issues associated with logging statements. Li et al. (Li et al., 2023a) focused on the readability of log messages in the logging statements, they derived three aspects related to the readability and explored the possibility of automatically classifying the readability of log messages. Ding et al. (Ding et al., 2023) made the first attempt to comprehensively study the temporal relations between logging statements and corresponding source codes. They proposed a tool to detect the issues of temporal inconsistencies. Bouzenia et al. (Bouzenia and Pradel, 2023) targeted checking for inconsistencies between conditions and logging statements and presented CMI-Finder, a neural network-based model to detect condition-statement inconsistencies. Chen et al. (Chen and Jiang, 2019, 2017) characterized and detected the anti-patterns in the logging statements. They encoded six anti-patterns into a static code analysis tool to detect previously unknown anti-patterns in the source code.
The primary distinction between our work and the aforementioned studies lies in the fact that while these studies solely detected specific categories that may affect the quality of logging statements, our approach not only identifies these issues but also provides solutions for their rectification.
Empirical studies on the existing logging practices. Owing to advancements in logging within software engineering, there has been a surge in research exploring logging practices. Li et al. (Li et al., 2021b) conducted an in-depth qualitative analysis of both the benefits and limitations of logging in software development. Zhou et al. (Zhou et al., 2020) delved into the ramifications of logging practices on data leaks within mobile apps. Kabinna et al. (Kabinna et al., 2018) identified that aspects like bug corrections, feature augmentations, and code restructuring often led to modifications in logging statements. Recently, Li et al. (Li et al., 2023b) pioneered a study on LLMs for logging statement generation and discovered that employing prompt-based strategies with zero-shot or few-shot learning enhances the generalization capabilities of LLMs.
8. Conclusion
In this paper, we first conduct a pilot study and identify four types of logging statement defects that developers are concerned about, which can affect the quality of runtime logs. Then we propose LogFixer, a framework for automatically detecting and updating potential defective logging statements. Experiment results demonstrate the effective detection and updating capabilities and practicality on new projects.